OpenStudy (anonymous):

Manuel traveled 4 hours nonstop to Mexico, a total of 172 miles. He took a train part of the way, which averaged 50mph, and then took a bus the remaining distance, which averaged 30mph. How long was Manuel on the train (in hours)?

OpenStudy (ybarrap):

average speed to mexico = 172 mi/4 hr train speed = 50 mi/hr bus speed = 30 mi/hr 50m/hr * tt + 30mi/hr * tb = 172 hrs, where tt = time on train, tb = time on bus tt50 + tb30 = 172 miles tb + tt = 4 hrs Ans. tt = 13/5 hrs and tb = 7/5 hrs

OpenStudy (anonymous):

Here is slightly different way to look at it. train_dist = train_rate * train_time bus_dist = bus_rate * bus_time train_dist + bus_dist = 172 Therefore: (train_rate * train_time) + (bus_rate * bus_time) = 172 Let x = train_time, then (4-x) = bus_time. Now substitute in the variables and the rates and you will have: (50 * x) + (30 * (4-x) ) = 172 50x + 120 -30x = 172 - 120 -120 ----------------------- 20x = 52 x = (52/20) = 13 / 5 = 2.6 hours

OpenStudy (anonymous):

r t = d Solve the following for t, the time on the train: 50 t + (4 - t) 30 = 172 t = 13/5 hours or 2hours and 36 minutes